OpenStudy (anonymous):

HELP MEDAL AND FAN!!!!!!! Multiply and simplify. 5 1/4 x 4/5 A.1/5 B.5 C.1/4 D.4

OpenStudy (anonymous):

hold on one sec

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

5?

OpenStudy (anonymous):

hold on i got it

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i'm pretty sure it's 5 because 4/5 of 5 is 4

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

i have another question

OpenStudy (anonymous):

5 1/4 = 5.25 0.8 = 4/5 5.25 x 0.8 = 4.2 so yeah i'm pretty sure it is 5 :)

OpenStudy (anonymous):

no tht is wrong

OpenStudy (anonymous):

Multiply and simplify. 5 1/4 ·8/9 · 8/21 A.80/189 B.1 7/9 C.5 16/189 D.5 445/756

OpenStudy (anonymous):

@kkay_light @MREAR53

OpenStudy (anonymous):

@Evil-cow-8 @itsbribro @ohlookidontunderstandanotherthing

OpenStudy (anonymous):

c